Suppose that a researcher, using data on class size (CS) and average test scores from 100 third-grade classes, estimates the OLS regression: 1. Test-Score = 520.4-5.82 × CS (20.4) (2.21) A classroom has 22 students. What is the regressions prediction for that classrooms average test score? Last year a classroom had 19 students, and this year it has 23 students. What is the regressions prediction for the change in average test score? The sample average class size across the 100 classrooms is 21.4. What is the sample average of the test scores across the 100 classrooms? Construct a 95% confidence interval for A, the regression slope coefficient Calculate the test statistic for the two-sided test of the null hypothesis Ho: A- 0. Do you reject the null hypothesis at the 5% level? At the 1% level? Calculate the test statistic for the two-sided test of the null hypothesis Ho: -5.6. Without doing any additional calculations, determine whether -5.6 is contained in the 95% confidence interval for A . Construct the 99% confidence interval for Ao . a. b. c. d. e. f. g.
Answer #1

a)
y^= 520.4 - 5.82 * 22
= 392.36


b)
change in average test score = -5.82 ( 23 - 19)
= -23.28

it will reduce by 23.28 points

c)
y^= 520.4 - 5.82 * 21.4
= 395.852

d)

df =n-2 = 98
t = 1.9845

(-5.82 - 1.9845 * 2.21 , -5.82 + 1.9845 * 2.21 )
= ( -10.2057 , -1.4343 )

e)
since 0 is not present in 95% confidence interval
we reject the null at 5 % level

I have solved first 4 parts by HomeworkLib policy

Post e) to g) parts again